# RISCORE: Enhancing In-Context Riddle Solving in Language Models through Context-Reconstructed Example Augmentation

### v3

@@ Abstract
Riddle-solving requires advanced reasoning skills, pushing LLMs to engage in abstract thinking and creative problem-solving, often revealing limitations in their cognitive abilities. In this paper, we examine the riddle-solving capabilities of LLMs using a multiple-choice format, exploring how different prompting techniques impact performance on riddles that demand diverse reasoning skills. To enhance results, we introduce RISCORE (RIddle Solving with COntext REcontruciton) a novel fully automated prompting method that generates and utilizes contextually reconstructed sentence-based puzzles in conjunction with the original examples to create few-shot exemplars. Our experiments demonstrate that RISCORE significantly improves the performance of language models in both vertical and lateral thinking tasks, surpassing traditional exemplar selection strategies across a variety of few-shot settings  [^1] Code for our experiments is available at: https://github.com/GiannisPana/RISCORE.
@@ 1 Introduction
Reasoning in Natural Language Processing has been a field of increasing interest, especially since the surge of Large Language Models (LLMs) which showcase several reasoning shortcomings Yu et al. (2024); Qiao et al. (2023). The logical gap between humans and LLMs’ advanced reasoning is further exposed when probing the puzzle solving capabilities of such models Giadikiaroglou et al. (2024), denoting the necessity of intricate datasets Lin et al. (2021); Szomiu and Groza (2021); Zhang and Wan (2022) that test and enhance complex model reasoning. In terms of advanced reasoning skills, lateral thinking datasets such as BrainTeaser Jiang et al. (2023b) stress the creative reasoning capabilities of models, requiring them to defy obvious logical associations and approach a more abstract way of thinking.
At the same time, prompting is a widespread technique of extracting reasoning patterns from LLMs Qiao et al. (2023). Especially Chain-of-Thought (CoT) prompting Wei et al. (2023), as naturally designed to trigger the extraction of reasoning chains, is indeed proven to be one successful and viable prompting strategy in practice, even though its functionality is only applicable on larger models. Other promising techniques, such as in-context learning, are highly capable when the right data samples are utilized as exemplars, but fail to properly drive LLMs otherwise Dong et al. (2024). While similarity-driven exemplar selection serves as a well-established rule of thumb Liu et al. (2022); Qin et al. (2024); Dong et al. (2024), the notion of similarity becomes more obscure when concerning reasoning patterns: for example, the sentences A man shaves everyday, yet keeps his beard long and Tom attends class every day but doesn’t do any homework Jiang et al. (2023b) in fact represent the exact same reasoning pattern, despite being rather distant semantically.
Driven by this observation, we assume that by “stripping” a sentence from linguistic features, while focusing on pure reasoning patterns prevalent in sentences, we can provide the model with the critical information to unlock the desired underlying reasoning pattern. Sentences that represent the same reasoning pattern are provided as “context reconstruction” counterparts in BrainTeaser data Jiang et al. (2023b) or can be automatically generated using the provided method. Therefore, incorporating such reconstructed instances as few-shot exemplars, together with their original counterparts could serve as an alternative to existing few-shot strategies that respects reasoning similarity and analogy. An example of the method is provided in Figure 1. To this end, we contribute to the following:
@@
  • We propose RISCORE, a novel prompting method designed to enhance the in-context riddle-solving capabilities of LLMs. As a supplement to this, we introduce an algorithm for generating contextually reconstructed multiple-choice riddles.
@@
Figure 1: Shows an example of few-shot (FS) prompting alongside the same method using a context-reconstructed example. The text in green represents the context reconstruction of Example 1, produced by the proposed method. The example in red was removed simply to maintain the same number of shots.
@@ 2 Related work
Multiple reasoning senses have been studied over the years using language models, including commonsense Sap et al. (2020), arithmetic Luo et al. (2024), abductive Zhao et al. (2023), inductive Han et al. (2024), deductive Sanyal et al. (2022), analogical reasoning Sultan and Shahaf (2022) and others. Linear thinking processes that exploit rules and logic, termed as “vertical thinking”, have been previously explored within popular datasets Lin et al. (2021); Bisk et al. (2020), unveiling interesting reasoning patterns of language models. On the other hand, creative thinking has been widely underexplored and often purposefully excluded from reasoning benchmarks Speer et al. (2017); Sap et al. (2019), leading to a significant gap in literature, especially given the emergent capabilities that larger models present Wei et al. (2022). Puzzle solving closely lies to creative reasoning Giadikiaroglou et al. (2024); Lin et al. (2021); Zhang and Wan (2022), since out-of-the-box thinking is mainly required. Going one step further, overriding default presuppositions and associations occurring when reasoning leads to “lateral thinking” processes, leading to more tricky puzzles as demonstrated in the BrainTeaser dataset Jiang et al. (2023b) for the first time. In our work, we focus on probing vertical and lateral puzzle-solving reasoning abilities of language models.
@@
Discovering reasoning patterns in LLMs is often performed via various types of prompting Qiao et al. (2023). Zero-shot strategies harness simple but intuitive instructions, such as “Let’s think step-by-step” that successfully improve LLM reasoning Kojima et al. (2023); however, the search space of such magic prompts seems indefinite. Querying for intermediate reasoning steps was framed in the seminal Chain-of-Thought (CoT) prompting Wei et al. (2023), proving that it is sufficient to elicit reasoning in larger models. Nevertheless, the placement of demonstrations in few-shot prompting faces several challenges Dong et al. (2024); specifically, the way of selecting these exemplars themselves is prone to instabilities Lu et al. (2022); Min et al. (2022), proving the nonnegotiable need for developing optimal exemplar selection techniques. Retrieval based on similarity is established as a default selection technique Liu et al. (2022); Qin et al. (2024); Dong et al. (2024), with related enhancement optimizing ordering of selected exemplars Wu et al. (2023). With a focus on reasoning tasks, complexity of reasoning paths Fu et al. (2023) or diversity of exemplars Zhang et al. (2022) have already proven their merits, shifting the weight towards uncovering the patterns hidden in data themselves that are able to advance reasoning, rather than similarity-driven placement based on semantics. Inspired from the above, we craft exemplars that promote the hidden reasoning patterns rather than the semantics of related data samples, maintaining simple similarity-based retrieval for their placement; by outperforming other prompting techniques, we demonstrate that highlighting those reasoning patterns is adequate for advanced LLM reasoning, without further few-shot engineering.
@@ 3.1 RISCORE prompting
Consider the following two riddles: R1: “A man shaves every day, yet keeps his beard long” and R2: “What has a beard but never needs to shave?”. While these riddles are semantically similar, follow the same structure, and refer to the same objects, their reasoning processes differ. In R1, where the answer is “A barber,” the word “beard” is used literally. In R2, where the answer is “A tree,” “beard” is used metaphorically, referring to the “beard” of certain trees, such as the “oak”. Thus, when using these riddles as context for a new riddle—such as “I plant seeds every day, yet don’t have a single plot”—the answer will depend on the interpretation of the phrase “plant seeds.” A metaphorical interpretation might lead to “A teacher,” while a literal one might suggest “A hydroponicist.” This illustrates how riddle context can significantly shape the reasoning process required for solving it.
In this example, however, instead of using R2, a contextually reconstructed version of R1R3: “Tom attends class every day but doesn’t do any homework”—would provide a clearer reasoning process to the model. Although R2 is semantically closer to R1, its metaphorical use of “beard” contrasts with the literal meaning in the original, potentially obscuring the reasoning pathway that the model should follow for solving a new riddle.
Building on the previous example, we propose the RISCORE prompting method, designed to enhance the in-context riddle-solving abilities of LLMs. The RISCORE method supplements each exemplar in FS learning with a contextually reconstructed version of itself. This preserves the original reasoning process while only altering its context. This approach enables the model to delineate a clear and coherent reasoning trajectory, which it can follow to effectively solve new riddles.
As illustrated in Figure 1, our approach builds on top of existing few-shot (FS) methods Dong et al. (2024); Wang et al. (2024); Sultan and Shahaf (2022), with RISCORE not involved in the FS selection process. The goal of our method is to augment FS samples with automatically generated context-reconstructed examples (detailed in Section 3.2). The addition of context reconstructed examples has, in most cases, proven to be more beneficial for the models’ performance, surpassing even the utility of real examples extracted from the dataset (see Section G).
@@ 3.2 Method for generating context reconstructions
In this section, we describe our approach for generating high-quality reconstructed Question-Answer pairs, which can be used as input to a model, alongside their counterparts, to enhance performance on lateral thinking problems. This process draws inspiration from the semi-automated pipeline introduced in the BrainTeaser paper Jiang et al. (2023b), but we extend it by fully leveraging the capabilities of LLMs to automate the procedure.
@@
The first step involves generating one contextually reconstructed Question-Answer pair per selected instance, temporarily ignoring distractors. To achieve this, we provide the LLM with the riddle, the correct answer, and a system prompt that outlines the task. The prompt instructs the model to analyze the given Question-Answer pair, understand the riddle, identify the reasoning process that links the question to the answer, and then generate a similar riddle that follows the same reasoning process, along with a corresponding correct answer. Since the model is provided with both the riddle and the correct answer, understanding the riddle becomes easier than solving it independently Sultan and Shahaf (2022). To further refine this approach, we apply the method in both Zero-shot (ZS) and Few-shot (FS) settings, using preexisting pairs of original and contextually reconstructed questions from BrainTeaser. The prompts used in this process can be found in App. A.2. After generating the Question-Answer pair, a filtering procedure was conducted to ensure the quality of the instance and its alignment with the riddles in the dataset. This process involved applying rules concerning both the structure of the riddle and the answer. These rules are dataset-specific and can be adjusted to suit different datasets, ensuring adaptability across various contexts. Further analysis regarding this step can be found in App. A.1.
@@
This process involves generating incorrect answers to complement the multiple-choice options. While this task may initially appear straightforward, it presents several significant challenges. First, the distractors must not only match the number of original options but also be guaranteed to be incorrect when compared to the correct answer. Additionally, they should not deviate excessively from the correct answer, as significant divergence could undermine the validity and challenge of the riddle. It is also essential to account for cases where the correct answer is “None of the above”, as observed in the BrainTeaser dataset. The length of the answers also plays a crucial role in this process. For instance, BrainTeaser includes mostly answers with more than four or five words, while RiddleSense provides typically single-word answers. These variations in length introduce additional challenges in crafting effective distractors. To address this, we propose two distinct methods for generating distractors: one for long and another for short distractors, respectively.
@@
In this approach, the original distractors from the riddle are inputted into a LLM alongside the new Question generated in Step 1. For each distractor (processed one at a time, except for “None of the above”), the model is prompted to modify its context by integrating it with the new riddle’s context. Regardless of the quality of this integration, the distractor is guaranteed to be incorrect, though lower-quality incorporation may slightly reduce the distractor’s overall effectiveness. If the original riddle’s answer differs from “None of the above” but the new riddle’s answer is, an additional distractor must be generated. In such cases, we simply prompt the LLM to generate a new distractor based on the generated Question-Answer pair. Details of the aforementioned step can be found in App. B.1.
@@
In the final step of this pipeline, we use the Question-Answer pair, randomly shuffle the distractor set, select the required number of distractors, and place the correct answer in a randomly assigned position. More details regarding the Prompting process for our method can be found in App. C.
@@ 4 Experiments
We evaluate various models using different datasets with our method and compare their performance against multiple baselines.
@@ 4.1 Datasets
For our experiments, we utilized two different datasets, namely BrainTeaser and RiddleSense. A comprehensive analysis of the datasets used in our experiments can be found in App D.
We selected BrainTeaser primarily for two key reasons. First, BrainTeaser provides high-quality, manually crafted context reconstructions, which establish an upper bound on the model’s performance when these are incorporated into the input. This upper bound also serves as a benchmark for evaluating the quality of riddles generated by the automated method. Second, the dataset consists of riddles that test the lateral thinking abilities of the models, a reasoning process that has proven as a challenging cognitive process for LLMs Giadikiaroglou et al. (2024).
Additionally, experiments were conducted using the RiddleSense dataset, which emphasizes vertical thinking, to evaluate the method’s applicability across different types of riddles. Given that hand-crafted context reconstructions are not available for RiddleSense, the automated method was utilized exclusively in this set of experiments.

### v4

@@ Abstract
Riddle-solving requires advanced reasoning skills, pushing Large Language Models (LLMs) to engage in abstract thinking and creative problem-solving, often revealing limitations in their cognitive abilities. In this paper, we examine the riddle-solving capabilities of LLMs using a multiple-choice format, exploring how different prompting techniques impact performance on riddles that demand diverse reasoning skills. To enhance results, we introduce RISCORE (RIddle Solving with COntext REcontruciton), a novel fully automated prompting method that generates and utilizes contextually reconstructed sentence-based puzzles in conjunction with the original examples to create few-shot exemplars. Our experiments demonstrate that RISCORE significantly improves the performance of language models in both vertical and lateral thinking tasks, surpassing traditional exemplar selection strategies across a variety of few-shot settings  [^1] Code for our experiments is available at: https://github.com/GiannisPana/RISCORE.
@@ 1 Introduction
Reasoning in Natural Language Processing has been a field of increasing interest, especially since the surge of Large Language Models (LLMs), which showcase several reasoning shortcomings Yu et al. (2024); Qiao et al. (2023). The logical gap between humans and LLMs’ advanced reasoning is further exposed when probing the puzzle solving capabilities of such models Giadikiaroglou et al. (2024), denoting the necessity of intricate datasets Lin et al. (2021); Szomiu and Groza (2021); Zhang and Wan (2022) that test and enhance complex model reasoning. In terms of advanced reasoning skills, lateral thinking datasets such as BrainTeaser Jiang et al. (2023b) stress the creative reasoning capabilities of models, requiring them to defy obvious logical associations and approach a more abstract way of thinking.
Figure 1: Standard FS prompting vs RISCORE prompting: by encouraging reasoning-based selection of exemplars (riddle in green) in place of semantic similarity selection (riddle in red), the model unlocks the reasoning pattern, guided towards the correct answer.
At the same time, prompting is a widespread technique of extracting reasoning patterns from LLMs Qiao et al. (2023). Especially Chain-of-Thought (CoT) prompting Wei et al. (2023), as naturally designed to trigger the extraction of reasoning chains, is indeed proven to be one successful and viable prompting strategy in practice, even though its functionality is only applicable on larger models. Other promising techniques, such as in-context learning, are highly capable when the right data samples are utilized as exemplars, but fail to properly drive LLMs otherwise Dong et al. (2024). While similarity-driven exemplar selection serves as a well-established rule of thumb Liu et al. (2022); Qin et al. (2024); Dong et al. (2024), the notion of similarity becomes more obscure when concerning reasoning patterns: for example, the sentences “A man shaves everyday, yet keeps his beard long” and “Tom attends class every day but doesn’t do any homework” Jiang et al. (2023b) in fact represent the exact same reasoning pattern, despite being rather distant semantically.
Driven by this observation, we assume that by “stripping” a sentence from linguistic features, while focusing on pure reasoning patterns prevalent in sentences, we can provide the model with the critical information to unlock the desired underlying reasoning pattern. Sentences that represent the same reasoning pattern are provided as “context reconstruction” counterparts in BrainTeaser data Jiang et al. (2023b) or can be automatically generated using our proposed method. Therefore, incorporating such reconstructed instances as few-shot exemplars, together with their original counterparts, offers an alternative few-shot strategy that prioritizes reasoning similarity and analogy. An example of the method is provided in Figure 2. To this end, we contribute to the following:
@@
  • We propose RISCORE (Example provided in Figure 1), a novel prompting method designed to enhance the in-context riddle-solving capabilities of LLMs. As a supplement to this, we introduce an algorithm for generating contextually reconstructed multiple-choice riddles.
@@
Figure 2: An overview of RISCORE, where the reconstructed instances, along with their original counterparts, are incorporated as exemplars in the few-shot setting to enhance the model’s riddle solving ability.
@@ 2 Related work
Multiple reasoning senses have been studied over the years using language models, including commonsense Sap et al. (2020), arithmetic Luo et al. (2024), abductive Zhao et al. (2023), inductive Han et al. (2024), deductive Sanyal et al. (2022), analogical reasoning Sultan and Shahaf (2022) and others. Linear thinking processes that exploit rules and logic, termed as “vertical thinking”, have been previously explored within popular datasets Lin et al. (2021); Bisk et al. (2020), unveiling interesting reasoning patterns of language models. On the other hand, creative thinking has been widely underexplored and often purposefully excluded from reasoning benchmarks Speer et al. (2017); Sap et al. (2019), leading to a significant gap in literature, especially given the emergent capabilities that larger models present Wei et al. (2022). Puzzle solving closely lies to creative reasoning Giadikiaroglou et al. (2024); Lin et al. (2021); Zhang and Wan (2022), since out-of-the-box thinking is mainly required. Going one step further, overriding default presuppositions and associations occurring when reasoning leads to “lateral thinking” processes, leading to more tricky puzzles, as demonstrated in the BrainTeaser dataset Jiang et al. (2023b) for the first time. In our work, we focus on probing vertical and lateral puzzle-solving reasoning abilities of language models via prompting.
Figure 3: An overview of the automated method for generating a context-reconstructed riddle.
@@
Discovering reasoning patterns in LLMs is often performed via various types of prompting Qiao et al. (2023). Zero-shot strategies harness simple but intuitive instructions, such as “Let’s think step-by-step” that successfully improve LLM reasoning Kojima et al. (2023); however, the search space of such magic prompts seems indefinite. Querying for intermediate reasoning steps was framed in the seminal Chain-of-Thought (CoT) prompting Wei et al. (2023), proving that it is sufficient to elicit reasoning in larger models. Nevertheless, the placement of demonstrations in few-shot prompting faces several challenges Dong et al. (2024); specifically, the way of selecting these exemplars themselves is prone to instabilities Lu et al. (2022); Min et al. (2022), proving the non-negotiable need for developing optimal exemplar selection techniques. Similarity-based retrieval is established as a default selection technique Liu et al. (2022); Qin et al. (2024); Dong et al. (2024), with related enhancement optimizing the ordering of selected exemplars Wu et al. (2023). With a focus on reasoning tasks, complexity of reasoning paths Fu et al. (2023) or diversity of exemplars Zhang et al. (2022) have already proven their merits, shifting the weight toward uncovering the patterns hidden in the data themselves that are able to advance reasoning, rather than similarity-driven placement based on semantics. Inspired by the above, we craft exemplars that promote the hidden reasoning patterns rather than the semantics of related data samples, maintaining simple similarity-based retrieval for their placement; by outperforming other prompting techniques, we demonstrate that highlighting those reasoning patterns is adequate for advanced LLM reasoning, without further few-shot engineering.
@@ 3.1 RISCORE prompting
Consider the following two riddles: R1: “A man shaves every day, yet keeps his beard long” and R2: “What has a beard but never needs to shave?”. Although these riddles are semantically similar, follow the same structure, and refer to the same objects, their reasoning processes differ. In R1, where the answer is “A barber,” the word “beard” is used in the context of human grooming and personal appearance. In R2, where the answer is “A tree,” “beard” is used in the botanical or natural context, referring to the “beard” of certain trees, such as the “oak”. Thus, when using these riddles as context for a new riddle—such as “I plant seeds every day, yet don’t have a single plot”—the answer will depend on the interpretation of the phrase “plant seeds”, based on its contextual framing. In a creative context, the correct answer is "An author". Authors “plant seeds” of ideas through their writing but do not have a literal “plot”, unless it refers to the abstract concept of a story plot, which adds layers of interpretation. On the other hand, “A farmer” is not the correct answer because a farmer typically works with physical plots of land to plant seeds, which contradicts the phrase “yet don’t have a single plot.”
In this example, however, instead of using R2, a contextually reconstructed version of R1R3: “Tom attends class every day but doesn’t do any homework”—would provide a clearer reasoning process for the model. Although R2 is semantically closer to R1, its use of “beard” in a natural or botanical context introduces a different reasoning pathway, illustrating how the contextual framing of a riddle can clarify or obscure its intended answer.
Building on the previous example, we propose the RISCORE prompting method, designed to enhance the in-context riddle-solving abilities of LLMs. The RISCORE method supplements each exemplar in FS learning with a contextually reconstructed version of itself. This preserves the desired reasoning process while only altering its context. This approach enables the model to delineate a clear and coherent reasoning trajectory, which it can then follow to effectively solve new riddles.
As illustrated in Figure 2, our approach extends existing few-shot (FS) methods Dong et al. (2024); Wang et al. (2024); Sultan and Shahaf (2022), with RISCORE not involved in the exemplar selection process [^2] We utilize semantic similarity for optimal exemplar selection, but other methods from literature may be used.. Specifically, the goal of our method is to augment FS samples with automatically generated context-reconstructed examples (detailed in Section 3.2). The inclusion of these context-adapted examples has, in most cases, demonstrated greater impact in improving model performance, even outperforming the use of real examples extracted directly from the dataset (see Section G).
@@ 3.2 Method for generating context reconstructions
In this section, we present our approach for generating high-quality contextually reconstructed riddles to serve as few-shot exemplars for a model in a Multiple-Choice Question Answering (MQA) format. These exemplars, combined with the original pairs, aim to enhance the model’s performance on lateral and vertical thinking tasks. Our method builds on the semi-automated pipeline proposed in the BrainTeaser paper Jiang et al. (2023b), but we advance this framework by fully harnessing the capabilities of LLMs to automate the process. An overview of the automated method for generating a context-reconstructed riddle is provided in Figure 3.
@@
The first step involves generating one contextually reconstructed Question-Answer pair per selected instance, temporarily ignoring distractors. Distractors constitute the incorrect answer options presented alongside the correct answer in multiple-choice settings, designed to appear plausible and test the depth of understanding or reasoning of the model. To achieve this, we provide the LLM with the riddle, the correct answer, and a system prompt that outlines the task. The prompt instructs the model to analyze the given Question-Answer pair, understand the riddle, identify the reasoning process that links the question to the answer, and then generate a similar riddle that follows the same reasoning process, along with a corresponding correct answer. Since the model is provided with both the riddle and the correct answer, understanding the riddle becomes easier than solving it independently Sultan and Shahaf (2022). To further refine this approach, we apply the method in both Zero-shot (ZS) and Few-shot (FS) settings, using pre-existing pairs of original and contextually reconstructed questions from BrainTeaser. The prompts used in this process can be found in App. A.2. After generating the Question-Answer pair, a filtering procedure was conducted to ensure the quality of the instance and its alignment with the riddles in the dataset. This process involves applying rules concerning both the structure of the riddle and the answer. These rules are dataset-specific and can be adjusted to suit different datasets, ensuring adaptability across various contexts. Further analysis regarding this step can be found in App. A.1.
@@
This process involves generating incorrect answers to complement the multiple-choice options. While this task may initially appear straightforward, it presents several significant challenges. First, the distractors must not only match the number of original options, but also be guaranteed to be incorrect when compared to the correct answer. Additionally, they should not deviate excessively from the correct answer, as significant divergence could undermine the validity and challenge of the riddle. It is also essential to account for cases where the correct answer is “None of the above”, as observed in the BrainTeaser dataset, ensuring that the other three alternatives are wrong enough.
The length of the answers also plays a crucial role in this process. For instance, BrainTeaser includes mostly answers with more than four or five words, while RiddleSense provides typically single-word answers. These variations in length introduce additional challenges in crafting effective distractors. To address this, we propose two distinct methods for generating distractors: one for long and another for short distractors, respectively.
@@
In this approach, the original distractors of each riddle are given to an LLM alongside the new Question generated in Step 1. For each distractor (processed one at a time, except for “None of the above”), the model is prompted to modify its context by integrating it with the context of the new riddle. Regardless of the quality of this integration, the distractor is guaranteed to be incorrect, though lower-quality incorporation may slightly reduce its overall effectiveness. If the original riddle’s answer differs from "None of the above", but the new riddle’s answer matches it, an additional distractor must be generated. In such cases, we simply prompt the LLM to generate a new distractor based on the generated Question-Answer pair. Details of the aforementioned step can be found in App. B.1.
@@
In the final step of this pipeline, we use the Question-Answer pair, randomly shuffle the distractor set, select the required number of distractors, and place the correct answer in a randomly assigned position, so that a possible position bias is eliminated. More details regarding the Prompting process for our method can be found in App. C.
@@ 4 Experiments
We evaluate various LLMs using different datasets with our method and compare their performance against multiple baselines.
@@ 4.1 Datasets
For our experiments, we utilize two different datasets, namely BrainTeaser and RiddleSense, addressing vertical and lateral reasoning respectively. A comprehensive analysis of the datasets used in our experiments can be found in App D.
We select BrainTeaser primarily for two key reasons. First, BrainTeaser provides high-quality, manually crafted context reconstructions, which establish an upper bound on the model’s performance when these are incorporated into the input. This upper bound also serves as a benchmark for evaluating the quality of riddles generated by the automated method. Second, the dataset consists of riddles that test the lateral thinking abilities of the models, a reasoning process that has proven as a challenging cognitive process for LLMs Giadikiaroglou et al. (2024). Additionally, experiments using RiddleSense are conducted, emphasizing on the vertical thinking abilities of models, while verifying our method’s broad applicability across different types of riddles. Given that hand-crafted context reconstructions are not available for RiddleSense, the automated method was utilized without ground-truth reconstructed instances in this set of experiments.